OpenStudy (anonymous):

A piece of art is in the shape of an equilateral triangle with sides of 7 in. Find the area of the piece of art. Round your answer to the nearest tenth. Answer A. 21.2 in2 B. 42.4 in2 C. 17.3 in2 D. none of these

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

any idea

OpenStudy (anonymous):

@gitahimart

OpenStudy (anonymous):

do you know heron's formulae?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

given sides a,b,c then area =\[\sqrt{s(s-a)(s-b)(s-c)}\] where s = (a+b+c)/2 try it

OpenStudy (radar):